BIAS CCD Tests
Between 2002-Nov and 2003-Oct, CCD Tests were regularly performed for all six combinations of the BIAS CCD controller Amplifier and Gain readout modes.
Current Hardware Configuration
The current Hardware configuration is 2222222222 which has been installed since 2002-11-16:
- Amplifier B in BLOW mode. This is the DEFAULT mode. Unless you specifically set an alternative mode this is the one used.
- Amplifier B in ABLOW mode
- Amplifier B in BHIGH mode
- Amplifier B in ABHIGH mode
- Amplifier A in ALOW mode
- Amplifier A in ABLOW mode
- Amplifier A in AHIGH mode
- Amplifier A in ABHIGH mode
Previous Hardware configurations
Results for previous hardware configurations...
FEROS @ ESO-1.52m
CCD Test results from FEROS at the ESO-1.52m (1998 -- 2002-Sep-30) are kept here.
The FEROS CCD system is a "standard" ESO CCD and Cryostat system controlled by a Copenhagen University Observatory CCD Controller system. It consists of the following components:
- One 2048×4096 EEV CCD detector, 15x15µm pixels.
- One Contiuous Flow Cryostat.
- One CUO Pre-Amp.
- Two CUO CCD Controller boxes (each consisting of a crate, power supply and 8 interface cards (see below). The two controllers are designated #1 and #2. A given controller configuration is described by a ten digit integer with each integer either a 1 or a 2. The ten digits refer to:
- Power Supply
- Sequencer board
- Video board
- Shutter & Heater Controller
- Bias board
- Two IBM compatible PC interface boards (#1 & #2). The two boards have different IO ports 0x500 an 0x520 (#1 & #2 respectively).
The following is a history of changes/works done on the FEROS CCD since 2002-Oct-01, the day it was transfered to the MPG/ESO-2.20m
- 2002-10-01: CCD dewar taken by LED for maintenance (baking at low vacuum, and pumping).
- 2002-10-01: CCD #1 Controller Electronics (conf = 1111111111) delivered to MPG/ESO-2.20m.
- 2002-10-10: Video Board #1 and ??? #1 discovered to be faulty. Sent to CUO for repair. Initially Video board #2 and ??? #2 were installed in crate #1 (i.e. conf = 1111121112???)
- 2002-10-12???: Complete #2 controller installed (conf = 2222222222)
- 2002-10-14???: Noise developes on CCD see image... CCDTests strongly affected...
- 2002-11-16: Noise source traced to preamplifier. Extra capacitor installed by A.Pizarro. Noise source elliminated. CCDTest returns to "normal"...